It is graduation time in a small island nation. There are a total number of 17,688 graduating from college and are actively looking for work. The college graduates are not included in the current employment measures for the country, which are given in the table. Answer the related questions taking into account the college graduates and the information in the table. Number of employed Number of unemployed Total population over age 16 335,915 15,483 415,999 What is the size of the labor force? size of labor force: How many people are now out of the labor force? out of the labor force: What is the unemployment rate? Round the answer to two places after the decimal. unemployment rate: What is the total labor force participation rate? Round the answer to two places after the decimal. labor force participation rate:
Answer #1

Labour force is the sum of number of workers employed and number of workers unemployed = 335915 + 15483 + 17688 = 369086

Number of people out of labour force = 415999 - 369086 = 46913

Unemployment rate is equal to unemployed workers divided by labour force = 33171/369086 = 8.99%

Labour force participation rate is equals to labour force divided by total adult population = 369086/415999 = 88.72%